# Identifying multi-compartment Hodgkin-Huxley models with high-density extracellular voltage recordings

### v1

@@ Appendix E CA1 Pyramidal Neuron (C10 Model)
Figure E.1: Morphology of the C10 model. LM-t: lacunosum moleculare (thin), LM-M: lacunosum moleculare (medium), LM-T: lacunosum moleculare (thick), RadDist: radiatum distal, RadMed: radiatum medial, RadProx: radiatum proximal, OriProx: oriens proximal, OriDist: oriens distal.
@@
Moreover, unlike the canonical HH model, the gating kinetics vary across anatomical regions [Tomko et al., 2021]. Gating variables evolve according to first-order kinetics:
@@
The attenuation factor Na_att is defined per section based on Migliore et al. [1999] (Table 14).
@@
Reversal potentials follow Cutsuridis et al. [2010], with a small adjustment to $E_{\mathrm{L}}$ to enforce a resting potential near $-70$ mV across compartments (Table 15).
@@
The maximum conductances for each region were taken directly from Cutsuridis et al. [2010] (Table 16).
@@
We used the following physiological constants: axial resistivity $R_{a}=150\ \Omega\cdot\mathrm{cm}$, specific membrane capacitance $C_{m}=1\ \mu\mathrm{F}/\mathrm{cm}^{2}$, and extracellular resistivity $\rho_{\mathrm{ext}}=300\ \Omega\cdot\mathrm{cm}$. While we assume homogeneous extracellular resistivity, this simplification neglects known spatial variability [Gold et al., 2006], which could be addressed in future work. Note that Cutsuridis et al. [2010] also consider additional channels, which we neglected here. Including them in our model could also be the subject of future experiments.
Synthetic extracellular recordings were generated by injecting a constant current of 1.0 nA into the soma for 30 ms. Voltage traces were simulated using a synthetic Neuropixel Ultra probe at a 40 kHz sampling rate, with noise standard deviation $\sigma_{obs}=1\mu\mathrm{V}$.
@@
We fit this model using standard full covariance EKF. All voltages and channel states were initialized at their steady-state values at rest, and thus we assumed the initial state to be known. We set the initial covariance matrix to $\Sigma_{0}=0.1I$. Observation noise was also assumed to be known, as commercial devices that record extracellular potentials typically provide noise specifications [Ye et al., 2024, Multi Channel Systems MCS GmbH, 2020]. Since the state dynamics were well-specified and deterministic, we fixed the standard deviation of the membrane potentials and channel states to small values: $\sigma_{v}=0.001\,\mathrm{mV}/\mathrm{ms}^{1/2}$, and $\sigma_{\text{state}}=0.0001\ \mathrm{ms}^{-1/2}$ (same for all the channels) for all compartments. These values ensured numerical stability while avoiding degenerate gradients during optimization.
We optimized maximum conductances and geometric parameters using the Adam optimizer. We tried multiple learning rates and initialized the maximum conductances uniformly at random within the bounds shown in Table 17. The rotation angles were initialized uniformly on $[0,2\pi)$, and translations were initialized to place the soma near its true location. We chose this initialization to reduce the risk of converging to a mirror-reflected solution, which produces identical extracellular signals when using a planar probe (see Section C). We ran 100 random initializations with different seeds and selected the result that achieved the highest marginal log-likelihood.
@@
Although the marginal log-likelihood converged to its true value after a few hundred iterations (fig. E.3F), and both the voltage traces (fig. E.3A–E) and compartment positions (fig. 5A) were accurately recovered across the cell, some maximum conductances did not converge to their true values or even continued to fluctuate (fig. E.2) —reflecting the well-known degeneracy in conductance-based neuron models, where multiple parameter combinations can produce nearly identical voltage dynamics.
@@
Figure E.2: Inference of the maximum conductances of the C10 model in all the regions. A: Soma, B: Axon, C: OriProx, D: OriDist, E: RadProx, F: RadMed, G: RadDist, H: LM. While several maximum conductances converged to their true values, many were still fluctuating.
@@
Figure E.3: Recovery of the membrane voltage in different regions of the pyramidal cell. A: Soma, B: Axon, C: RadDist, D: OriDist, E: LM. F: Marginal log-likelihood trace during optimization. Despite some maximum conductances failing to converge to their true values, the membrane voltage was recovered accurately across the cell and the marginal log-likelihood converged to its true value, reflecting the well-known degeneracy in conductance-based neuron models.
Appendix F DiagEKF vs. EKF
@@
We compare the computational efficiency of the standard Extended Kalman Filter (EKF) and our sparse diagonal approximation (DiagEKF). To isolate performance differences, we consider a simplified neuron model: a $24\,\mu\mathrm{m}$ cable equipped with standard Hodgkin–Huxley channels [Hodgkin and Huxley, 1952]. A constant current is injected for 20 ms, and extracellular voltage is recorded at a single electrode at a 40 kHz sampling rate.
@@
The neuron is discretized into $N$ compartments, with $N\in\{1,10,20,50,100,200\}$. As $N$ increases, the number of state variables (membrane voltages and gating variables) grows proportionally. Importantly, the gating dynamics in these models are highly local: each variable typically depends only on its own previous value and the voltage of its corresponding compartment. As a result, the Jacobian of the dynamics function quickly becomes very sparse when the number of states increases (fig. F.1A).
@@
To evaluate performance, we ran both EKF and DiagEKF for 500 iterations on an NVIDIA H100 GPU for each value of $N$. As shown in fig. F.1B, DiagEKF achieves approximately 2× speedup over the dense EKF in high-dimensional settings. This highlights the computational advantage of our diagonal approximation in realistic biophysical neuron models.
In our work, we used DiagEKF to fit extracellular voltage recordings simulated from a retinal ganglion cell. All the other experiments were performed using standard EKF.
@@
Figure F.1: Computational efficiency of our diagonal approximation to the EKF filtering distribution A: Sparsity of the Jacobian of the dynamics function over state dimension. Especially large biophysical models are sparse, suggesting that sparse approximations to the filtering covariance could provide accurate but efficient inference. B: Average runtime over $500$ updates for the EKF and sparse diagEKF versus state dimension for a $20ms$ time horizon. The EKF with diagonal approximation is twice as fast as EKF with a dense covariance for large models.

### v2

@@ Appendix G CA1 Pyramidal Neuron (C10 Model)
Figure G.1: Morphology of the C10 model. LM-t: lacunosum moleculare (thin), LM-M: lacunosum moleculare (medium), LM-T: lacunosum moleculare (thick), RadDist: radiatum distal, RadMed: radiatum medial, RadProx: radiatum proximal, OriProx: oriens proximal, OriDist: oriens distal.
@@
Moreover, unlike the canonical HH model, the gating kinetics vary across anatomical regions [tomko2021new]. Gating variables evolve according to first-order kinetics:
@@
The attenuation factor Na_att is defined per section based on migliore1999role (Table 14).
@@
Reversal potentials follow cutsuridis2010encoding, with a small adjustment to $E_{\mathrm{L}}$ to enforce a resting potential near 70-70 mV across compartments (Table 15).
@@
The maximum conductances for each region were taken directly from cutsuridis2010encoding (Table 16).
@@
We used the following physiological constants: axial resistivity $R_{a}=150\ \Omega\cdot\mathrm{cm}$, specific membrane capacitance $C_{m}=1\ \mu\mathrm{F}/\mathrm{cm}^{2}$, and extracellular resistivity $\rho_{\mathrm{ext}}=300\ \Omega\cdot\mathrm{cm}$. While we assume homogeneous extracellular resistivity, this simplification neglects known spatial variability [gold2006origin], which could be addressed in future work. Note that cutsuridis2010encoding also consider additional channels, which we neglected here. Including them in our model could also be the subject of future experiments.
Synthetic extracellular recordings were generated by injecting a constant current into the soma for 30 ms. Voltage traces were simulated using a synthetic Neuropixel Ultra probe at a 40 kHz sampling rate, with noise standard deviation $\sigma_{obs}=0.1\mu\mathrm{V}$.
@@
We fit this model using the block-diagonal EKF. All voltages and channel states were initialized at their steady-state values at rest, and thus we assumed the initial state to be known. We set the initial covariance matrix to $\mathbf{\Sigma}_{0}=0.1\mathbf{I}$. Observation noise was also assumed to be known, as commercial devices that record extracellular potentials typically provide noise specifications [ye2024ultra, MEA2100manual]. Since the state dynamics were well-specified and deterministic, we fixed the standard deviation of the membrane potentials and channel states to small values: $\sigma_{v}=0.001\,\mathrm{mV}/\mathrm{ms}^{1/2}$, and $\sigma_{\text{state}}=0.0001\ \mathrm{ms}^{-1/2}$ (same for all the channels) for all compartments. These values ensured numerical stability while avoiding degenerate gradients during optimization.
We optimized maximum conductances and geometric parameters using the Adam optimizer. We tried multiple learning rates and initialized the maximum conductances uniformly at random within the bounds shown in Table 17. The rotation angles were initialized uniformly on $0,2\pi)$, and translations were initialized to place the soma near its true location. We chose this initialization to reduce the risk of converging to a mirror-reflected solution, which produces identical extracellular signals when using a planar probe (see Section D). We ran 100 random initializations with different seeds and selected the result that achieved the highest marginal log-likelihood.
@@
Although the marginal log-likelihood converged to its true value after a few hundred iterations (fig.˜G.3F), and both the voltage traces (fig.˜G.3A–E) and compartment positions (fig.˜5A) were accurately recovered across the cell, the leak conductances did not converge to their true values or even continued to fluctuate (fig.˜G.2) —reflecting the well-known degeneracy in conductance-based neuron models, where multiple parameter combinations can produce nearly identical voltage dynamics.
@@
Figure G.2: Inference of the maximum conductances of the C10 model in all the regions. A: Soma, B: Axon, C: OriProx, D: OriDist, E: RadProx, F: RadMed, G: RadDist, H: LM. While most maximum conductances converged to their true values, many–especially the leak conductances–were still fluctuating.
@@
Figure G.3: Recovery of the membrane voltage in different regions of the pyramidal cell. A: Soma, B: Axon, C: RadDist, D: OriDist, E: LM. F: Marginal log-likelihood trace during optimization. Despite some maximum conductances failing to converge to their true values, the membrane voltage was recovered accurately across the cell and the marginal log-likelihood converged to its true value, reflecting the well-known degeneracy in conductance-based neuron models.
Appendix H DiagEKF vs. EKF runtime
@@
We compare the computational efficiency of the standard Extended Kalman Filter (EKF) and our sparse diagonal approximation (DiagEKF). To isolate performance differences, we consider a simplified neuron model: a $24\,\mu\mathrm{m}$ cable equipped with standard Hodgkin–Huxley channels [hodgkin1952quantitative]. A constant current is injected for 20 ms, and extracellular voltage is recorded at a single electrode at a 40 kHz sampling rate.
@@
The neuron is discretized into $N$ compartments, with $N\in\{1,10,20,50,100,200\}$. As $N$ increases, the number of state variables (membrane voltages and gating variables) grows proportionally. Importantly, the gating dynamics in these models are highly local: each variable typically depends only on its own previous value and the voltage of its corresponding compartment. As a result, the Jacobian of the dynamics function quickly becomes very sparse when the number of states increases (fig.˜H.1A).
@@
To evaluate performance, we ran both EKF and DiagEKF for 500 iterations on an NVIDIA H100 GPU for each value of $N$. As shown in fig.˜H.1B, DiagEKF achieves approximately 2× speedup over the dense EKF in high-dimensional settings. This highlights the computational advantage of our diagonal approximation in large biophysical neuron models.
@@
Figure H.1: Computational efficiency of our diagonal approximation to the EKF filtering distribution A: Sparsity of the Jacobian of the dynamics function over state dimension. Especially large biophysical models are sparse, suggesting that sparse approximations to the filtering covariance could provide accurate but efficient inference. B: Average runtime over $500$ updates for the EKF and sparse diagEKF versus state dimension for a $20ms$ time horizon. The EKF with diagonal approximation is twice as fast as EKF with a dense covariance for large models.